# Relativistic Guiding-Center Motion: Action Principle, Kinetic Theory, and Hydrodynamics

### v1

@@ Abstract
Further work is needed to incorporate collisions into the guiding-center kinetic equation. In real astrophysical problems, one needs to have in mind that in addition to the electrons there are also ions. From the kinetic equation with collisions term one can find out the structure of the dissipative terms in the hydrodynamic equations, and to compute the kinetic coefficients [26]. Some of these coefficients are also quantities of great interest in two-dimensional physics: Hall conductivity, odd viscosity [27, 28], thermoelectric [29] and thermal Hall coefficient. To find out the behavior of these coefficients in quantizing magnetic fields, one needs to understand the collisions between guiding centers.
Finally, in an electromagnetic plasma, the guiding-center kinetic and hydrodynamic equations are coupled to Maxwell equations. We defer a full treatment of the resulting theories to future work.
@@
The authors are indebted to Jingyuan Chen, Masaru Hongo, Sanjay Reddy, and Eliot Quataert for discussions. The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University and RIKEN iTHEMS, where part of this work was completed during the workshop (YITP-T-23-05) on “Condensed Matter Physics of QCD 2024.” This work is supported, in part, by the U.S. DOE Grants Nos. DE-FG02-13ER41958 and DE-FG02-01ER41195 and by the Simons Collaboration on Ultra-Quantum Matter, which is a Grant from the Simons Foundation (No. 651440, DTS).
@@
The Lagriangian (2) is a quadratic polynomial of $p$. Integrating out $p$, i.e., using Eq. (5a), we find
@@
Compared to the Lagrangian of a particle, the guiding-center Lagrangians (S4) and (S5) involve an additional projector $\tilde{\Delta}_{\mu\nu}$. The projector makes sure that only the motion along the magnetic field carries the usual kinetic energy, while the energy of the fast synchrotron motion is included in the rest mass $\tilde{m}$ of the guiding-center “particle.”
@@
with arbitrary parameter $v_{\parallel}$. Substituting Eqs. (S7) and keeping only the leading term in $\epsilon_{*}$ in each of the three (almost) orthogonal directions $\bm{B}$, $\bm{E}\times\bm{B}$ and $\bm{E}$ we find:
@@
The first term is the (arbitrary) velocity along the magnetic field direction. The second term is the usual drift velocity in the crossed electric and magnetic fields:
@@
Acoording to Eq. (5a), total drift in the direction perpendicular to $\bm{B}$ is given by the sum of $\bm{v}_{\perp}$ and $\bm{v}_{\perp}^{*}$, which occur even if the fields are uniform, and $\bm{v}_{\text{D}}$ which occurs only in non-uniform or time-dependent fields. The latter drift is given by the spatial components of vector $v_{\text{D}}^{\mu}$ in Eq. (6):
@@
Following the standard practice we shall introduce a unit vector in the direction of the magnetic field $\bm{\bm{b}}\equiv\bm{B}/B$. Using $\bm{E}=\bm{B}\times\bm{v}_{\perp}$ one can show that every term in Eq. (S12) is perpendicular to $\bm{b}$. In particular,
@@
where we used Eq. (S9) and neglected $\mathcal{O}(\epsilon_{*})$ term. Putting together Eqs. (S12), (S13), (S10) and (S11) we obtain for the total drift velocity perpendicular to the magnetic field the following expression:
@@
This result coincides with Eq. (243) in Ref. [13] in the form presented in Eq. (1.76) of Ref. [14] with identification $\bm{b}\to\bm{e}_{1}$, $\bm{v}_{\perp}\to\bm{u}_{E}$ and $B_{*}\to\sqrt{B^{2}-E_{\perp}^{2}}$. Note that $\tilde{m}\tilde{\gamma}$ in Eq. (S14) is the total energy of the particle (including the energy of its motion around the guiding center), which in Ref. [14] is denoted by $m_{0}\gamma$.
@@
where the Liouville operator $\mathcal{L}$ was defined in Eq. (7), and $\dot{x}$ and $\dot{p}$ are given in Eqs. (5) and (6), taken to be valid in the whole 8-dimensional phase space. We search for a solution with support on the intersection of the mass-shell hyperboloid $p^{2}=\tilde{m}^{2}$ and the hyperplane $B^{\mu\nu}p_{\nu}=0$. We also require that particles have positive energy, so
@@
The constraints $p^{2}-\tilde{m}^{2}=0$ and $B^{\mu\nu}p_{\nu}=0$ are maintained by the equation of motion: $\mathcal{L}[p^{2}-\tilde{m}^{2}]=\mathcal{L}[B^{\mu\nu}p_{\nu}]=0$. In order to find the action of the Liouville operator on $\delta(p^{2}-\tilde{m}^{2})\delta^{2}(B^{\mu\nu}p_{\nu})$, one needs to find how the arguments of these delta function evolve outside the constraint hypersurface. We have
@@
where $\tilde{c}$ is a numerical coefficient. To find $\tilde{c}$ it is sufficient to consider the case of a constant magnetic field pointing along the $z$ direction. Choosing the affine parameter along the particle trajectory to be $x^{0}$, we have, from Eq. (5a), $\alpha=1/p^{0}$ and thus, substituting Eqs.(S16) and (S23) into Eq. (9),
@@
Therefore, $\tilde{c}=1/(2\pi^{2})$.
@@
$W=\frac{\tilde{c}}{\alpha}\left(B_{*}+B_{*}^{-1}p_{\mu}\partial_{\nu}B^{\mu\nu}\right)\theta(p^{0})\delta(p^{2}-m^{2})\delta^{2}(p_{i})=\frac{\tilde{c}}{\alpha}\left(B_{*}^{3}+B_{*}p_{\mu}\partial_{\nu}B^{\mu\nu}\right)\theta(p^{0})\delta(p^{2}-m^{2})\delta^{2}(B^{\mu\nu}p_{\nu})\,.$ (S42)

### v2

@@ Abstract
Further work is needed to incorporate collisions into the guiding-center kinetic equation. In real astrophysical problems, one needs to have in mind that in addition to the electrons, there are also ions. From the kinetic equation with collisions term, one can determine the structure of the dissipative terms in the hydrodynamic equations and compute the kinetic coefficients [39]. Some of these coefficients are also quantities of great interest in two-dimensional physics: Hall conductivity, odd viscosity [40, 41], thermoelectric [42], and thermal Hall coefficient. To determine the behavior of these coefficients in quantizing magnetic fields, one needs to understand the collisions between guiding centers.
Finally, our system of relativistically invariant equations for the guiding-center motion can be coupled to Maxwell fields via the full electromagnetic current given in Eq. (25), which includes the drift current (21) and the magnetization current. These ingredients, together with Eq. (10) for the invariant phase-space volume, should open a new route for numerical simulations of relativistic astrophysical plasmas using the particle-in-cell or other methods [43, 44].
@@
Acknowledgement—The authors are indebted to Jingyuan Chen, Masaru Hongo, Sanjay Reddy, Anatoly Spitkovsky, and Eliot Quataert for discussions. The authors thank the Yukawa Institute for Theoretical Physics at Kyoto University and RIKEN iTHEMS, where part of this work was completed during the workshop (YITP-T-23-05) on “Condensed Matter Physics of QCD 2024.” This work is supported, in part, by the U.S. DOE Grants No. DE-FG02-13ER41958 and No. DE-FG02-01ER41195, by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (No. 651440, D.T.S.), and by the United-States-Israel Binational Science Foundation (BSF) (Grant No. 2022110).
Appendix: Guiding-center dynamics in curved spacetime—In curved spacetime, Eq. (2) becomes
$L=-p_{\mu}\dot{x}^{\mu}-A_{\mu}\dot{x}^{\mu}+\frac{\alpha}{2}(g^{\mu\nu}p_{\mu}p_{\nu}-\tilde{m}^{2})+\lambda_{\alpha}B^{\alpha\nu}p_{\nu}.$ (A1)
The equations of motion following from this Lagrangian are the general coordinate invariant generalization of Eqs. (5):
$\dot{x}^{\mu}$ $=\alpha g^{\mu\nu}p_{\nu}+\lambda_{\alpha}B^{\alpha\mu}\equiv\alpha p^{\mu}+v_{\mathrm{D}}^{\mu},$ (A2a)
$\dot{p}_{\mu}-\dot{x}^{\nu}\Gamma^{\alpha}_{\nu\mu}p_{\alpha}$ $=F_{\mu\nu}\dot{x}^{\nu}+\frac{\alpha}{2}J\nabla_{\mu}B_{*}-\lambda_{\alpha}p_{\beta}\nabla_{\!\mu}B^{\alpha\beta}.$ (A2b)
In addition, one has two constraints $p^{2}-\tilde{m}^{2}=0$ and $B^{\alpha\nu}p_{\nu}=0$. Substituting Eq. (A2a) into Eq. (A2b) and following the same steps that led to Eq. (6), we find
$v^{\mu}_{\mathrm{D}}=\alpha B_{*}^{-2}\!\left(p_{\nu}(p\cdot\nabla)B^{\mu\nu}+\frac{J}{2}B^{\mu\nu}\nabla_{\nu}B_{*}\right).$ (A3)
For a nonrelativistic particle in a constant magnetic field and a weak gravitational field ($g_{00}=1+2\Phi$, $\Phi\ll 1$), Eq. (A3) gives the standard gravitational drift velocity $-mB^{-2}\bm{\nabla}\Phi\times\bm{B}$ [46]. It is interesting to note that one and the same term in Eq. (A3) is responsible for the magnetic field curvature drift and for the gravitational drift. This is a manifestation of Einstein’s equivalence principle. For example, consider a particle moving along a circular magnetic field line. In the frame rotating with the particle there is a centripetal force which causes the same gravitational drift as the curvature drift in the original frame. This argument is a rigorous version of a textbook derivation of the curvature drift [10].
The invariant phase-space volume is given by
$W=\frac{B_{*}+B_{*}^{-1}p_{\mu}\nabla_{\nu}B^{\mu\nu}}{2\pi^{2}\alpha}\theta(p_{0})\delta(p^{2}-\tilde{m}^{2})\delta^{2}(e_{i}^{\mu}p_{\mu}),$ (A4)
where $e_{i}^{\mu}$, $i=1,2$, are two orthonormal basis vectors in the plane perpendicular to the magnetic field: $e_{i}^{\mu}e_{i}^{\nu}=\Delta^{\mu\nu}$, $g_{\mu\nu}e_{i}^{\mu}e^{\nu}_{j}=\delta_{ij}$.
[…]
— Supplemental Material —
Relativistic Guiding-Center Motion: Action Principle, Kinetic Theory, and Hydrodynamics
Dam Thanh Son and Mikhail Stephanov
@@
The Lagrangian (2) is a quadratic polynomial of $p$. Integrating out $p$, i.e., using Eq. (5a), we find
@@
Compared to the Lagrangian of a particle, the guiding-center Lagrangians (S4) and (S5) involve an additional projector $\tilde{\Delta}_{\mu\nu}$. The projector makes sure that only the motion along the magnetic field carries the usual kinetic energy, while the energy of the fast cyclotron motion is included in the rest mass $\tilde{m}$ of the guiding-center “particle.”
@@
with arbitrary parameter $v_{\parallel}$. Substituting Eqs. (S7) and keeping only the leading term in $\epsilon_{*}$ in each of the three (almost) orthogonal directions $\bm{B}$, $\bm{E}\times\bm{B}$ and $\bm{E}$, we find
@@
The first term is the (arbitrary) velocity along the magnetic field direction. The second term is the usual drift velocity in crossed electric and magnetic fields:
@@
According to Eq. (5a), total drift in the direction perpendicular to $\bm{B}$ is given by the sum of $\bm{v}_{\perp}$ and $\bm{v}_{\perp}^{*}$, which occur even if the fields are uniform, and $\bm{v}_{\text{D}}$ which occurs only in non-uniform or time-dependent fields. The latter drift is given by the spatial components of vector $v_{\text{D}}^{\mu}$ in Eq. (6):
@@
Following the standard practice, we shall introduce a unit vector in the direction of the magnetic field $\bm{\bm{b}}\equiv\bm{B}/B$. Using $\bm{E}=\bm{B}\times\bm{v}_{\perp}$, one can show that every term in Eq. (S12) is perpendicular to $\bm{b}$. In particular,
@@
where we used Eq. (S9) and neglected $\mathcal{O}(\epsilon_{*})$ term. Putting together Eqs. (S12), (S13), (S10), and (S11), we obtain, for the total drift velocity perpendicular to the magnetic field, the following expression:
@@
This result coincides with Eq. (243) in Ref. [18] in the form presented in Eq. (1.76) of Ref. [19] with the identification $\bm{b}\to\bm{e}_{1}$, $\bm{v}_{\perp}\to\bm{u}_{E}$ and $B_{*}\to\sqrt{B^{2}-E_{\perp}^{2}}$. Note that $\tilde{m}\tilde{\gamma}$ in Eq. (S14) is the total energy of the particle (including the energy of its motion around the guiding center), which in Ref. [19] is denoted by $m_{0}\gamma$.
@@
where the Liouville operator $\mathcal{L}$ was defined in Eq. (7), and $\dot{x}$ and $\dot{p}$ are given in Eqs. (5) and (6), taken to be valid in the whole 8-dimensional phase space. We search for a solution with support on the intersection of the positive energy $p^{0}>0$ mass-shell hyperboloid $p^{2}=\tilde{m}^{2}$ and the hyperplane $B^{\mu\nu}p_{\nu}=0$, i.e.,
@@
The constraints $p^{2}-\tilde{m}^{2}=0$ and $B^{\mu\nu}p_{\nu}=0$ are maintained by the equations of motion: $\mathcal{L}[p^{2}-\tilde{m}^{2}]=\mathcal{L}[B^{\mu\nu}p_{\nu}]=0$. In order to find the action of the Liouville operator on $\delta(p^{2}-\tilde{m}^{2})\delta^{2}(B^{\mu\nu}p_{\nu})$, one needs to find how the arguments of these delta function evolve outside the constraint hypersurface. We have
@@
where $\tilde{c}$ is a numerical coefficient.
To find $\tilde{c}$ it is sufficient to consider the case of a constant magnetic field pointing along the $z$ direction. Choosing the affine parameter along the particle trajectory to be $x^{0}$, we have, from Eq. (5a), $\alpha=1/p^{0}$ and thus, substituting Eqs.(S16) and (S23) into Eq. (9),
@@
Therefore, $\tilde{c}=1/(2\pi^{2})$. Substituting that into Eq. (S23), and using $\delta^{2}(B^{\mu\nu}p_{\nu})=B_{*}^{-2}\delta^{2}(\Delta^{\mu\nu}p_{\nu})$ in Eq. (S16), we obtain Eq. (10) up to $\mathcal{O}(\partial)$ terms.
@@
$W=\frac{\tilde{c}}{\alpha}\left(B_{*}+B_{*}^{-1}p_{\mu}\partial_{\nu}B^{\mu\nu}\right)\theta(p^{0})\delta(p^{2}-m^{2})\delta^{2}(p_{i})=\frac{\tilde{c}}{\alpha}\left(B_{*}+B_{*}^{-1}p_{\mu}\partial_{\nu}B^{\mu\nu}\right)\theta(p^{0})\delta(p^{2}-m^{2})\delta^{2}(\Delta^{\mu\nu}p_{\nu}).$ (S42)